OpenStudy (ashy98):

help me please solve the equation??? 0.5x-1.4=1.1

OpenStudy (usukidoll):

add 1.4 to both sides and divide by 0.5

undeadknight26 (undeadknight26):

like so:D .5x - 1.4 = 1.1 + 1.4 +1.4 .5x = 2.5 /.5 /5 x =?

OpenStudy (usukidoll):